### v2

# Efficient Semi-Supervised Adversarial Training via Latent Clustering-Based Data Reduction

Abstract
Achieving high model robustness under adversarial settings is widely recognized as demanding considerable training samples. Recent works propose semi-supervised adversarial training (SSAT) methods with external unlabeled or synthetically generated data, which are the current state-of-the-art. However, SSAT requires substantial extra data to attain high robustness, resulting in prolonged training time and increased memory usage. In this paper, we propose unlabeled data reduction strategies to improve the efficiency of SSAT. Specifically, we design novel latent clustering-based techniques to select or generate a small critical subset of data samples near the model’s decision boundary. While focusing on boundary-adjacent points, our methods maintain a balanced ratio between boundary and non-boundary data points to avoid overfitting. Comprehensive experiments on benchmark datasets demonstrate that our methods can significantly reduce SSAT’s data requirement and computation costs while preserving its strong robustness advantages. In particular, our latent-space selection scheme based on k-means clustering and our guided DDPM fine-tuning approach with LCG-KM are the most effective, achieving nearly identical robust accuracies with $5\times$ to $10\times$ less unlabeled data and approximately $4\times$ less total runtime.
I Introduction
Deep neural networks (DNNs) are highly vulnerable to adversarially perturbed inputs [1, 2]. This phenomenon has raised serious concerns about the reliability of DNNs in safety-critical applications and has driven numerous research into designing methods to enhance model robustness [2, 3, 4, 5]. Among them, adversarial-training-based algorithms are most popular [7, 8, 9]. As illustrated by Schmidt et al. [10], adversarially robust learning is inherently challenging, requiring a much larger sample complexity than standard learning. To address this, semi-supervised adversarial training (SSAT) has been proposed to enlarge the training set size with either external unlabeled data [11, 12] or generated synthetic data [13, 14, 15]. Despite alleviating the sample complexity barrier and producing models with improved robustness performance, these SSAT algorithms typically utilize vast amounts of additional data, suggesting the requirement of much larger hardware to store the extra data and a much longer training time for algorithms like adversarial training to converge (see Appendix A for detailed discussions on related work).
Motivated by the above observations, we propose to study whether the large amount of extra unlabeled data is inevitable for SSAT to achieve high robust accuracy. The ultimate goal is to maximize model robustness while utilizing as few additional data samples as possible. Inspired by Zhang et al. [16], which highlights the unequal importance of training examples, we argue that with limited model capacity, SSAT should also focus on critical points near the model’s decision boundary. Specifically, we introduce three data selection strategies, including PCS, a simple approach that prioritizes unlabeled data points with low prediction confidence, and two advanced latent clustering-based selection schemes: LCS-KM and LCS-GMM, where we identify the most uncertain data in the model’s latent embedding space based on the corresponding clustering technique. The selected data points are then incorporated into the SSAT pipeline as the unlabeled dataset. In addition, we propose novel DDPM fine-tuning approaches with guidance loss inspired by our data selection schemes, including PCG, LCG-KM, and LCG-GMM, to further reduce the computation overhead of SSAT algorithms that require the pre-generation of synthetic data using diffusion models. Our data reduction schemes balance the ratio of boundary critical points and the remaining non-boundary points to avoid potential overfitting to the boundary distribution. By applying the aforementioned data reduction schemes, the high unlabeled data and computation requirements of SSAT can be significantly alleviated while largely preserving its robustness performance.
Contributions. We motivate and formalize the problem tasks of reducing the volume of unlabeled data while maintaining model robustness for SSAT (Section II). To realize this goal, we propose to optimize the performance of SSAT by refining the model’s decision boundary in the regions of high uncertainty by strategically prioritizing boundary-adjacent unlabeled data points (Section III-A). Specifically, we propose various data selection schemes to identify boundary-adjacent points (Section III-B), including a simple method based on prediction confidence and two advanced selection techniques by clustering the unlabeled or generated data in the latent embedding space. Additionally, we introduce three guided DDPM fine-tuning approaches (Section III-C), aiming to generate the subset of desirable data directly, thus avoiding the computation overhead of pre-generating the whole extra dataset.
Comprehensive experiments on image benchmarks demonstrate that our data selection schemes, especially for the best-performing LCS-KM approach, significantly improve the data and computational efficiency of baseline SSAT methods, while achieving comparable robust accuracies (Section IV-A). For instance, utilizing $10\%$ extra unlabeled data identified by LCS-KM in SSAT can achieve similar robustness performance on SVHN compared to using the entire dataset, while reducing the total runtime to approximately $25\%$. By directly generating the exact amount of boundary-adjacent data points, our guided DDPM fine-tuning approach, such as LCG-KM, can reduce the computation overhead while maintaining robustness (Section IV-C). Moreover, we demonstrate the generalizability of our data reduction schemes to a real-world COVID-related medical application (Section V-A) and visualize the distributions of selected data points to further illustrate why our LCS-KM method always performs the best (Section V-B). Finally, we provide general guidelines on how to choose the key hyperparameters for maximizing the SSAT performance (Section V-C), and conclude with more discussions on related literature, limitations of our study, and future research directions (Section VI). Our work reveals the importance of prioritizing boundary-adjacent points for improving SSAT’s efficiency and highlights the superiority of latent-based k-means clustering, potentially assisting the development of more efficient and scalable robust learning algorithms for real-world applications.
II Inefficiency of Existing SSAT Methods
II-A Semi-Supervised Adversarial Training
We work with adversarially robust classification tasks and focus on semi-supervised adversarial training (SSAT), which was initially introduced to the field by Carmon et al. [11]. Let $\mathcal{X}\subseteq\mathbb{R}^{d}$ be the input space and $\mathcal{Y}$ be the output space of class labels. Assume $D_{l}$ is the underlying labeled data distribution over $\mathcal{X}\times\mathcal{Y}$ that we aim to learn adversarially robust classifiers. Let $\mathcal{S}_{l}$ be the training dataset with examples independently and identically drawn from $D_{l}$, and $\mathcal{S}_{u}$ be a collection of inputs sampled from an unlabeled data distribution $D_{u}$ defined over $\mathcal{X}$. In adversarial ML literature, $D_{u}$ is typically considered to be similar but not identical to the marginal input distribution of $D_{l}$, and the set size of $\mathcal{S}_{u}$, denoted as $|\mathcal{S}_{u}|$, to be much larger than $|\mathcal{S}_{l}|$, since unlabeled data are much easier to acquire than well-annotated labeled data. For instance, when $\mathcal{S}_{l}$ corresponds to the 5050K labeled CIFAR-10 training images, $\mathcal{S}_{u}$ is the 500500K “most CIFAR-10-like but non-identical” images selected from the whole 8080M Tiny ImageNet dataset [11].
Specifically, SSAT usually first trains an intermediate model $f_{\hat{\theta}}$ on the labeled training set $\mathcal{S}_{l}$, known as the pseudo labeler, and then assigns pseudo-labels to each unlabeled data point $\bm{x}\in\mathcal{S}_{u}$. The weight parameters of the intermediate model are optimized by minimizing the following objective function:
$\hat{\theta}=\operatorname*{argmin}_{\theta}\ \frac{1}{|\mathcal{S}_{l}|}\sum_{(\bm{x},y)\in\mathcal{S}_{l}}\ell(\theta,\bm{x},y),$ (1)
where $\ell(\cdot)$ stands for the individual standard loss (e.g., cross-entropy), capturing the discrepancy between model predictions and the ground-truth labels. Finally, SSAT adversarially trains a classification model, denoted as $\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{S}_{u},\gamma)$, on both labeled data $\mathcal{S}_{l}$ and unlabeled data $\mathcal{S}_{u}$ but paired with pseudo labels by minimizing the following adversarial loss:
$\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{S}_{u},$ $\gamma)=\operatorname*{argmin}_{\theta}\quad\frac{\gamma}{|\mathcal{S}_{l}|}\sum_{(\bm{x},y)\in\mathcal{S}_{l}}\ell_{\mathrm{adv}}(\theta,\bm{x},y)$
$+\frac{1-\gamma}{|\mathcal{S}_{u}|}\sum_{\bm{x}\in\mathcal{S}_{u}}\ell_{\mathrm{adv}}\!\left(\theta,\bm{x},f_{\hat{\theta}}(\bm{x})\right),$ (2)
where $\gamma\geq 0$ is a hyperparameter that controls the trade-off between labeled and unlabeled data. A typical approach used in prior work is to construct each training batch with varying ratios of labeled and unlabeled data (corresponding to different values of $\gamma$) but fix the batch size to optimize for the best SSAT performance [13]. For simplicity, we term $\gamma$ as the per-batch ratio. When $\gamma$ is set as $1$, Equation II-A reduces to the optimization objective of vanilla adversarial training (AT) [7]. In Equation II-A, $\ell_{\mathrm{adv}}(\cdot)$ denotes the individual adversarial loss. Specifically, for any $(\bm{x},y)\in\mathcal{S}_{l}$ and strength $\epsilon>0$,
$\ell_{\mathrm{adv}}(\theta,\bm{x},y)=\max_{\bm{\delta}\in\mathcal{B}_{\epsilon}(\bm{0})}\ell(\theta,\bm{x}+\bm{\delta},y),$ (3)
where $\mathcal{B}_{\epsilon}(\bm{0})$ denotes the perturbation ball centered at $\bm{0}$ with radius $\epsilon$ in some distance metric such as $\ell_{p}$-norm. Aligned with prior literature [7], we adopt multi-step PGD to approximately solve for the optimization problem in Equation 3.
II-B Limitations of Prior Work
Data Inefficiency. Although incorporating unlabeled data alleviates the sample complexity barriers of adversarially robust learning [10] and can produce models with higher robust accuracies, they require a substantial amount of extra unlabeled data $\mathcal{S}_{u}$ to ensure effective robustness enhancement. Such a trend is clearly documented in the leaderboard of RobustBench [17]. For CIFAR-10 image classification tasks, state-of-the-art SSAT algorithms either select 500500K external Tiny Imagenet samples or generate millions of synthetic CIFAR-10-like images, both significantly larger than the original 5050K CIFAR-10 training images. Specifically, the method proposed by Gowal et al. [13] achieves around $65\%$ robust accuracy on CIFAR-10 using WideResNet-70-16 against $\ell_{\infty}$ perturbations with radius $\epsilon=8/255$, increasing the robustness performance of vanilla AT [7] by a large margin; however, it relies on an extra unlabeled set of 100100M DDPM-generated data. Training SSAT algorithms on the huge unlabeled dataset demands significantly more computing power, such as specialized hardware, data parallelism, or distributed training, which in turn induces much higher energy consumption and carbon footprint. In addition, handling a larger dataset in ML often means increased GPU processing load and memory usage.
High Computation Cost. Apart from the data inefficiency issue, SSAT also requires a longer convergence time to obtain the best-performing model, usually $2$ to $4$ times the number of total training epochs compared with vanilla AT (see Figure 1 for supporting evidence). Intuitively, learning meaningful and robust representations through adversarial training from such a large and potentially diverse unlabeled dataset is more difficult. As our experiments illustrate, the slower convergence of SSAT can be attributed mainly to the large set of unlabeled data (paired with pseudo labels), which often induces higher gradient variance than the original labeled data samples. Once we reduce the size of the unlabeled data while retaining the essential information, SSAT algorithms are observed to converge much faster. We note that vanilla AT has already been criticized for its high computation [18, 19], due to the iterative PGD steps involved in dealing with the inner maximization problem (Equation 3). SSAT methods incur even higher computations, since they converge more slowly and usually adopt larger model architectures (see Table I for concrete runtime comparisons). Therefore, developing more efficient methods, especially for resource-constrained scenarios, becomes critical for SSAT to effectively learn the core information in the unlabeled dataset.
II-C Problem Formulation
Witnessing the data and computational inefficiency of SSAT methods, we propose to study whether the large set size of unlabeled data is inevitable for achieving high model robustness. To be more specific, we introduce two general methodology frameworks for improving the efficiency of SSAT algorithms, corresponding to the following optimization problems:
Strategic Selection. Inspired by the idea of coreset selection for data-efficient deep learning [20, 21, 22], the first approach aims to strategically search for a small but essential set of unlabeled data $\mathcal{A}_{u}\subseteq\mathcal{S}_{u}$ such that SSAT with the selected subset $\mathcal{A}_{u}$ can produce models with comparable robustness to those obtained using the full unlabeled dataset $\mathcal{S}_{u}$. Formally, we aim to solve the following constrained optimization problem:
$\max_{\mathcal{A}_{u}\subseteq\mathcal{S}_{u}}\mathrm{AdvRob}_{\epsilon}\big(\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{A}_{u},\gamma)\big),\text{ s.t. }|\mathcal{A}_{u}|\leq\alpha|\mathcal{S}_{u}|,$ (4)
where $\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{A}_{u},\gamma)$ denotes the model learned by SSAT with $\mathcal{S}_{l}$ and $\mathcal{A}_{u}$ based on Equation II-A, and $\alpha\in(0,1)$ is a predefined ratio capturing the data constraint. Here, $\mathrm{AdvRob}_{\epsilon}(\theta)$ denotes the model robustness on $\mathcal{D}_{l}$ under $\epsilon$ perturbations:
$\mathrm{AdvRob}_{\epsilon}(\theta)\!=\!1\!-\!\mathbb{E}_{(\bm{x},y)\sim D_{l}}\!\left[\max_{\bm{\delta}\in\mathcal{B}_{\epsilon}(\bm{0})}\!\ell_{0/1}(\theta,\bm{x}+\bm{\delta},y)\right]$ (5)
where $\ell_{0/1}$ denotes the 0-1 loss function. Due to the combinatorial nature and the high computation costs of AT algorithms, it is computationally hard to enumerate all the feasible subsets $\mathcal{A}_{u}$ to solve the proposed optimization problem exactly. As we will illustrate in Section III-B, we design different selection schemes for the extra unlabeled dataset $\mathcal{S}_{u}$ that are effective in approximately solving the optimization problem in Equation 4 while only incurring negligible computational overhead.
Guided Diffusion. The second method proposes to leverage existing pretrained diffusion models but fine-tune them to directly generate a small set of extra data with desirable properties, ready to be incorporated into the SSAT pipeline. Similar to our first strategic selection approach, we aim to achieve comparable robustness to existing SSAT while maximizing the efficiency for the utilization of the additional data:
$\max_{\mathcal{G}_{u}}\mathrm{AdvRob}_{\epsilon}\big(\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{G}_{u},\gamma)\big),\text{ s.t. }|\mathcal{G}_{u}|\leq\alpha|\mathcal{S}_{u}|,$ (6)
where $\mathcal{G}_{u}$ denotes a small set of unlabeled data generated by a fine-tuned generative model. As will be detailed in Section III-C, we adopt a pretrained DDPM [23] as the backbone while strategically fine-tuning it with novel training losses for guided data generations. Compared with the previous selection-based method, a guided diffusion approach can avoid the inefficient pre-generation of a large amount of unlabeled data, making it particularly suitable for SSAT methods that already utilize a pre-trained diffusion model for synthetic data generation.
III Improving the Efficiency of SSAT via Latent Clustering-Based Data Reduction
So far, we have formulated the problem tasks of reducing the unlabeled dataset size to improve SSAT’s efficiency while maximizing the robustness enhancement. In this section, we first explain why boundary-adjacent data points are essential for more efficient utilization of unlabeled data (Section III-A), then we detail the proposed method designs for both strategic selection and guided generation (Sections III-B and III-C).
III-A Prioritize Boundary-Adjacent Unlabeled Data
Achieving high robustness with restricted data resources remains challenging, as it involves striking a balance between data significance, optimizing the use of unlabeled data, and tackling constraints imposed by model capacity. Inspired by prior literature [16, 24] that emphasizes the unbalanced data importance for vanilla adversarial training (see Appendix A for more discussions), we hypothesize that not all unlabeled data contribute equally to the robustness enhancement for SSAT. In particular, we propose to identify or generate a small set of vulnerable yet valuable unlabeled data points, which are close to the model’s decision boundary. Consequently, such boundary-adjacent data points are highly susceptible to label changes under small input perturbations and are inherently difficult for the model to classify. Thus, improving their classification can yield more robustness enhancement to adversarial inputs. We expect the decision boundary of the intermediate model $f_{\hat{\theta}}$ to act as an effective proxy function for locating difficult-to-classify data points. Since the intermediate model is usually trained with strong standard accuracy, it is expected to preserve class semantics and provide a solid foundation for identifying these critical points. By focusing more on these critical boundary points with respect to the intermediate model, the final model can be trained more efficiently using SSAT algorithms while upholding robust accuracies.
While a plausible approach is to select a set of unlabeled data based on how much their label predictions change under adversarial perturbations returned by PGD attacks, it is computationally expensive, undermining the efficiency gains we aim for. In particular, such a PGD-based method entails iterative optimization processes to pinpoint boundary points, making them impractical for large-scale unlabeled datasets. Therefore, we seek better alternatives to effectively identify boundary data points without incurring high computational overhead.
Algorithm 1 Prediction Confidence-Based Selection
1:  Input: labeled $\mathcal{S}_{l}$; unlabeled $\mathcal{S}_{u}$; selection ratio $\alpha$; perturbation size $\epsilon$; boundary ratio $\beta$; per-batch ratio $\gamma$
2:  $\hat{\theta}\leftarrow$ train a standard classification model on $\mathcal{S}_{l}$
3:  $\hat{y}\leftarrow$ predict pseudo label $f_{\hat{\theta}}(\bm{x})$ for $\bm{x}\in\mathcal{S}_{u}$
4:  Initialize $\mathcal{A}_{u}$ as empty set
5:  $\mathrm{Conf}(\bm{x})\leftarrow$ get confidence of $f_{\hat{\theta}}(\bm{x})$ for $\bm{x}\in\mathcal{S}_{u}$
6:  Sort $\mathcal{S}_{u}$ by ascending order of $\mathrm{Conf}(\bm{x})$
7:  $\mathcal{A}_{u}\leftarrow$ add top $\beta\cdot\alpha|\mathcal{S}_{u}|$ points with lowest $\mathrm{Conf}(\bm{x})$
8:  $\mathcal{A}_{u}\leftarrow$ add $(1-\beta)\cdot\alpha|\mathcal{S}_{u}|$ points randomly from $\mathcal{S}_{u}\setminus\mathcal{A}_{u}$
9:  ${\theta}_{\mathrm{final}}\leftarrow$ $\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{A}_{u},\gamma)$
10:  Output: selected subset $\mathcal{A}_{u}$, final model ${\theta}_{\mathrm{final}}$
III-B Strategic Selection via Latent Clustering
PCS. To identify a small, critical subset of boundary-adjacent data points while accounting for computational efficiency and scalability, we first propose a straightforward approach, coined prediction confidence-based selection (PCS), which makes use of the intermediate model $f_{\hat{\theta}}$ to compute a prediction confidence score for each unlabeled data point. The pseudocode for such a selection scheme is depicted in Algorithm 1. Initially, all the data points in $\mathcal{S}_{u}$ are sorted by their prediction confidence $\mathrm{Conf}(\cdot)$, with those exhibiting the lowest confidence scores being prioritized. The underlying assumption is that data points with low confidence scores are more likely to lie near the model’s decision boundary, making them ideal candidates for our selection. Note that the parameter $\beta\in[0,1]$ is introduced to balance the ratio between boundary and non-boundary points to avoid overfitting, which is used in all our proposed selection schemes. The reason for involving such a trade-off parameter will be further discussed in Section V-C.
The biggest advantage of PCS is its high computational efficiency for ranking the unlabeled data. Note that for SSAT, we usually have an order-wise larger collection of unlabeled data than the original labeled dataset. Thus, an efficient selection scheme is desirable to avoid high computational overhead. Nevertheless, our experiments showed that using prediction confidence may not capture the underlying structure of the data well, leading to decreased robustness enhancement when incorporated in SSAT. We hypothesize that PCS overlooks the geometric relationships and distributional properties, which can be crucial for characterizing boundary-adjacent points. In addition, DNNs have been shown to be overconfident in their predictions [25], suggesting that prediction confidence scores returned by neural network models might be inherently biased.
LCS. To overcome the aforementioned issues, we introduce latent clustering-based selection (LCS) schemes to pinpoint desirable data in the model’s latent space using different clustering techniques. Our approach begins by generating latent embeddings for all $\bm{z}=h_{\hat{\theta}}(\bm{x})$ with $\bm{x}\in\mathcal{S}_{u}$, where $h_{\hat{\theta}}$ denotes the mapping of the input layer to the penultimate layer with respect to the intermediate model $f_{\hat{\theta}}$. The penultimate layer captures more abstract, high-level features and better represents the underlying data structure, which helps identify points near decision boundaries. It avoids the biases of overconfident predictions from the final layer, offering more reliable clustering. Here, the goal is to identify boundary-adjacent data points inferred by examining distances to cluster centroids. Points equidistant from multiple centroids are expected to be closer to decision boundaries in the latent embedding space.
In particular, we explore two classical clustering techniques for their unique benefits in identifying boundary points in the LCS framework. Algorithm 2 presents the pseudocode of the two proposed LCS methods. In particular, LCS with k-means (LCS-KM) generates latent representations of unlabeled data and clusters them based on Euclidean distances to the centroids, prioritizing data points that are equidistant from multiple centroids to capture local geometric structures around decision boundaries effectively. On the other hand, LCS with Gaussian mixture models (LCS-GMM) fits the latent representations to Gaussian mixture models, using posterior probabilities across multiple fitted Gaussians to identify points that are likely near decision boundaries. Compared to PCS, both LCS-KM and LCS-GMM leverage latent space clustering to provide a more accurate characterization of boundary vulnerabilities.
Algorithm 2 Latent Clustering-Based Selection
1:  Input: labeled $\mathcal{S}_{l}$; unlabeled $\mathcal{S}_{u}$; selection ratio $\alpha$; cluster number $k$; perturbation size $\epsilon$; boundary ratio $\beta$; per-batch ratio $\gamma$
2:  $\hat{\theta}\leftarrow$ train a standard classification model on $\mathcal{S}_{l}$
3:  $\hat{y}\leftarrow$ predict pseudo label $f_{\hat{\theta}}(\bm{x})$ for $\bm{x}\in\mathcal{S}_{u}$
4:  Initialize $\mathcal{A}_{u}$ as empty set
5:  $\bm{z}\leftarrow$ get latent embeddings $h_{\hat{\theta}}(\bm{x})$ for $\bm{x}\in\mathcal{S}_{u}$
6:  if using LCS-KM then
7:   $\{\mathcal{C}_{1},\ldots,\mathcal{C}_{k}\}\leftarrow$ k-means clustering on $\{\bm{z}\}_{\bm{x}\in\mathcal{S}_{u}}$
8:   $\Delta d\leftarrow$ compute $\Delta d=|d_{1}-d_{2}|$ for $\bm{x}\in\mathcal{S}_{u}$, where $d_{1}$ and $d_{2}$ are Euclidean distances to the nearest two centroids
9:   $\mathcal{A}_{u}\leftarrow$ add top $\beta\cdot\alpha|\mathcal{S}_{u}|$ points with smallest $\Delta d$
10:  else if using LCS-GMM then
11:   Fit a GMM with $k$ components to $\{\bm{z}\}_{\bm{x}\in\mathcal{S}_{u}}$
12:   Compute GMM posterior probability $\bm{p}$ for each $\bm{z}$
13:   $\Delta p\leftarrow$ compute $\Delta p=|p_{1}-p_{2}|$ for each $\bm{z}$, where $p_{1}$ and $p_{2}$ are the top two highest probabilities
14:   $\mathcal{A}_{u}\leftarrow$ add top $\beta\cdot\alpha|\mathcal{S}_{u}|$ points with smallest $\Delta\gamma$
15:  end if
16:  $\mathcal{A}_{u}\leftarrow$ add $(1-\beta)\cdot\alpha|\mathcal{S}_{u}|$ points randomly from $\mathcal{S}_{u}\setminus\mathcal{A}_{u}$
17:  ${\theta}_{\mathrm{final}}\leftarrow$ $\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{A}_{u},\gamma)$
18:  Output: selected dataset $\mathcal{A}_{u}$, final model ${\theta}_{\mathrm{final}}$
LCS-KM. In this variant of LCS, we first generate latent embeddings $\{\bm{z}=h_{\hat{\theta}}(\bm{x}):\bm{x}\in\mathcal{S}_{u}\}$, and then partition the unlabeled dataset into $k$ clusters $\{\mathcal{C}_{1},\mathcal{C}_{2},\ldots,\mathcal{C}_{k}\}$ by minimizing the within-cluster sum of squares $\sum_{j=1}^{k}\sum_{\bm{z}\in\mathcal{C}_{j}}\|\bm{z}-\bm{\mu}_{j}\|^{2}$, where $\bm{\mu}_{j}$ is the centroid of the $j$-th cluster. For each $\bm{z}$, we compute the Euclidean distance to each cluster centroid $\|\bm{z}-\bm{\mu}_{j}\|$. Data points are selected based on the minimal difference in distance between each latent embedding $\bm{z}$ to the corresponding two closest cluster centroids $\Delta d=|d_{1}-d_{2}|$, where $d_{1}$ and $d_{2}$ are the Euclidean distances to the closest and second closest centroids, respectively. Finally, the top $\alpha|\mathcal{S}_{u}|$ points from $\mathcal{S}_{u}$ with the smallest $\Delta d$ values form the reduced unlabeled dataset, followed by the same balancing step using the ratio parameter $\beta$. As will be demonstrated in our experiments, prioritizing these strategically selected unlabeled data points during semi-supervised adversarial training can achieve comparable robustness with much-improved efficiency.
LCS-GMM. In this variant of LCS, we again start by computing latent embeddings for unlabeled data $\{\bm{z}=h_{\hat{\theta}}(\bm{x}):\bm{x}\in\mathcal{S}_{u}\}$. Instead of using k-means clustering, we fit these latent representations using Gaussian mixture models (GMMs). A GMM assumes that the data is generated from a mixture of $k$ Gaussian distributions, each with its own mean $\bm{\mu}_{j}$ and covariance matrix $\mathbf{\Sigma}_{j}$ for the $j$-th component. Mathematically, each data point $\bm{z}$ has a probability of belonging to the $j$-th Gaussian component given by the posterior probability:
$p_{j}(\bm{z})=\frac{\pi_{j}\cdot\mathcal{N}(\bm{z}\mid\bm{\mu}_{j},\mathbf{\Sigma}_{j})}{\sum_{i=1}^{k}\pi_{i}\cdot\mathcal{N}(\bm{z}\mid\bm{\mu}_{i},\mathbf{\Sigma}_{i})},$ (7)
where $\pi_{j}$ is the mixing coefficient for the $j$-th Gaussian component, and $\mathcal{N}(\bm{z}\mid\bm{\mu}_{j},\mathbf{\Sigma}_{j})$ denotes the Gaussian distribution with mean $\bm{\mu}_{j}$ and covariance $\Sigma_{j}$. We focus on data points with similar top posterior probabilities, suggesting that they are near the boundary between Gaussian distributions. Specifically, for each $\bm{z}$, we calculate the difference between the highest and the second-highest posterior probabilities: $\Delta p=|p_{1}(\bm{z})-p_{2}(\bm{z})|$, where $p_{1}(\bm{z})$ and $p_{2}(\bm{z})$ are the highest and second-highest posterior probabilities, respectively. Data points with the smallest $\Delta\gamma$ values are selected, as they are more likely to be near decision boundaries. Finally, the top $\alpha|\mathcal{S}_{u}|$ unlabeled data with the smallest $\Delta p$ values are used to form the selected subset. As will be illustrated in Section IV, the robustness performance of SSAT can be largely maintained when using LCS-KM and LCS-GMM to select a small subset of unlabeled data (e.g., $\alpha=10\%$ or $\alpha=20\%$). More detailed discussions of the difference between these data selection schemes are provided in Section V-B, where we visualize the selected unlabeled data in a two-dimensional latent space.
Algorithm 3 SSAT with Guided DDPM Fine-Tuning
1:  Input: labeled $\mathcal{S}_{l}$; pre-trained DDPM $\theta_{\mathrm{pre}}$; fine-tuning epochs $S$; trade-off parameter $\lambda$; other parameters $\epsilon$, $\alpha|\mathcal{S}_{u}|$, $\beta$, $\gamma$, $k$, $\eta$
2:  $\hat{\theta}\leftarrow$ train a standard classification model on $\mathcal{S}_{l}$
3:  Initialize $\mathcal{G}_{u}$ as empty set
4:  $\mathcal{G}_{u}\leftarrow$ add $(1-\beta)\cdot\alpha|\mathcal{S}_{u}|$ points generated by pre-trained $\theta_{\mathrm{pre}}$
5:  Initialize $\theta$ as $\theta_{\mathrm{pre}}$
6:  for $s=1,2,\ldots,S$ do
7:   Sample batches of size $b$: $\bm{x}_{0}\sim\mathcal{S}_{l}$, $t\sim\mathcal{U}(T)$, $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\mathbf{I})$
8:   $\bm{x}_{t}\leftarrow\sqrt{\bar{\alpha}_{t_{i}}}\bm{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}$
9:   $L_{\mathrm{DDPM}}\leftarrow\frac{1}{b}\sum_{i\in[b]}\|\bm{\epsilon}-\bm{\epsilon}_{\theta}(\bm{x}_{t},t)\|_{2}^{2}$
10:   if using PCG then
11:    $L_{\mathrm{reg}}\leftarrow\frac{1}{b}\sum_{i\in[b]}\|\ell_{\mathrm{PC}}(g_{\theta}(\bm{x}_{t},t))\|_{2}^{2}$
12:   else if using LCG-KM then
13:    $L_{\mathrm{reg}}\leftarrow\frac{1}{b}\sum_{i\in[b]}\|\ell_{\mathrm{KM}}(g_{\theta}(\bm{x}_{t},t))\|_{2}^{2}$
14:   else if using LCG-GMM then
15:    $L_{\mathrm{reg}}\leftarrow\frac{1}{b}\sum_{i\in[b]}\|\ell_{\mathrm{GMM}}(g_{\theta}(\bm{x}_{t},t))\|_{2}^{2}$
16:   end if
17:   $\theta\leftarrow\theta-\eta\nabla_{\theta}\big(L_{\mathrm{DDPM}}+\lambda\cdot L_{\mathrm{reg}}\big)$
18:  end for
19:  $\mathcal{G}_{u}\leftarrow$ add $\beta\cdot\alpha|\mathcal{S}_{u}|$ points generated by fine-tuned $\theta$
20:  ${\theta}_{\mathrm{final}}\leftarrow$ $\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{G}_{u},\gamma)$
21:  Output: generated dataset $\mathcal{G}_{u}$, final model $\theta_{\mathrm{final}}$
III-C Guided Diffusion via Latent Clustering
While our selection-based strategy can significantly reduce the unlabeled data set size and lower the overall training time, it has a critical drawback–when SSAT employs a generative approach, all our previous methods require generating the full synthetic unlabeled dataset as a prerequisite before selection, which is inefficient in terms of both generation time and storage requirement. As shown in Table II, generating 1M CIFAR-10-like synthetic images via DDPM requires around $4$ hours on a single A100 GPU. Note that the best current SSAT method documented in the RobustBench leaderboard [17] utilizes a few hundred million synthetically generated images, suggesting the total generation time can be a significant computational overhead. Therefore, a natural question arises: whether pre-generating synthetic data and then applying strategic selection schemes can be further improved. In this section, we propose an alternative generative pipeline, detailed in Algorithm 3, which answers the above question affirmatively. In particular, we fine-tune a pretrained DDPM using novel guidance losses for more efficient unlabeled data generation based on the insights from previous sections. In this way, we directly generate a small, critical set of boundary-adjacent data points, thereby avoiding redundant data generation and further reducing the computation overhead.
DDPM. Diffusion model resembles a milestone in the history of generative modeling, which is the current state-of-the-art for high-quality image generation and synthesis. A denoising diffusion probabilistic model (DDPM) is a deep neural network trained to learn how to step-by-step reverse the forward process of adding Gaussian noise to an image [23]. By gradually removing Gaussian noise through the reverse process, DDPM can thus generate real-looking images from pure noise. Let $q_{0}$ be a real image distribution over $\mathbb{R}^{d}$ and $\mathcal{N}(\bm{0},\mathbf{I})$ be the noise distribution. The forward process defines how to map images $\bm{x}_{0}\sim q_{0}$ to Gaussian noises $\bm{x}_{T}\sim\mathcal{N}(\bm{0},\mathbf{I})$ through conditional sampling: for any $t\in[T]$, let
$q(\bm{x}_{t}|\bm{x}_{t-1})=\mathcal{N}(\bm{x}_{t};\sqrt{1-\beta_{t}}\bm{x}_{t-1},\beta_{t}\mathbf{I}),$ (8)
where $\beta_{1},\beta_{2},\ldots,\beta_{T}$ are scheduling parameters. According to the properties of Gaussian distributions, Equation 8 implies that we can directly sample $\bm{x}_{t}$ for any $t\in[T]$ based on:
$q(\bm{x}_{t}|\bm{x}_{0})=\mathcal{N}(\bm{x}_{t};\sqrt{\bar{\alpha}_{t}}\bm{x}_{0},(1-\bar{\alpha}_{t})\mathbf{I}),$ (9)
where $\bar{\alpha}_{t}=\Pi_{s=1}^{t}(1-\beta_{s})$. Based on Equation 9, Ho et al. [23] proposed to train DDPM by minimizing a regression loss:
$L_{\mathrm{DDPM}}(\theta)=\mathbb{E}_{\bm{x}_{0}\sim q_{0},\bm{\epsilon}\sim\mathcal{N}(\bm{0},\mathbf{I}),t\in\mathcal{U}(T)}\big\|\bm{\epsilon}-\bm{\epsilon}_{\theta}(\bm{x}_{t},t)\big\|_{2}^{2},$ (10)
where $\bm{x}_{t}=\sqrt{\bar{\alpha}_{t}}\bm{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}$ denotes the noisy image corresponding to $\bm{x}_{0}$ at $t$-th forward step, $\bm{\epsilon}_{\theta}:\mathbb{R}^{d}\times[0,1]\rightarrow\mathbb{R}^{d}$ stands for a neural net with weights $\theta$ for noise prediction from $\bm{x}_{t}$, and $\mathcal{U}(T)$ denotes the uniform distribution over the set $\{1,2,\ldots,T\}$. Once the model $\bm{\epsilon}_{\theta}$ is trained, we can then generate synthetic images from Gaussian noise via the following reverse (or denoising) process: for $t=T,\ldots,1$,
$\bm{x}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\big(\bm{x}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\bm{\epsilon}_{\theta}(\bm{x}_{t},t)\big)+\sqrt{\beta_{t}}\bm{\delta},$ (11)
where $\alpha_{t}=1-\beta_{t}$, and $\bm{\delta}\sim\mathcal{N}(\bm{0},\mathbf{I})$. For simplicity, denote by $g_{\theta}(\bm{x}_{t},t)$ the right-hand side of Equation 11, and $g_{\theta}:\mathbb{R}^{d}\times[0,1]\rightarrow\mathbb{R}^{d}$ can be viewed as one-step DDPM denoising.
Define DDPM Guidance Loss. Now that we have explained DDPM’s forward and backward processes, we can introduce the detailed design of our fine-tuning schemes. Borrowing the insights from previous sections, our method aims to fine-tune a DDPM model to directly generate the required subset of boundary-adjacent unlabeled data points for more efficient SSAT. Similar to our selection-based methods, we consider three designs: prediction confidence-guided generation (PCG), latent clustering-guided generation with k-means (LCG-KM), and latent clustering-guided generation with Gaussian mixture models (LCG-GMM). Specifically, we define a guidance loss $\ell_{\mathrm{guide}}:\mathbb{R}^{d}\rightarrow\mathbb{R}_{+}$ corresponding to each design, which will serve as a fine-tuning regularization term to bias the diffusion model to generate more boundary-adjacent samples.
PCG. The first design utilizes the prediction confidence of the intermediate model $f_{\hat{\theta}}$ and defines the guidance loss as
[…]
A larger value of $\ell_{\mathrm{PC}}(\bm{x})$ suggests that $\bm{x}$ is more likely to be closer to the model’s decision boundary. As we’ve explained in Section III-B, the prediction confidence of a neural network classifier can be an unreliable estimator. Therefore, we propose alternative guidance losses using more robust latent clustering-based approaches. Based on different clustering techniques, we again introduce two possible variants.
LCG-KM. For LCG-KM, we first obtain the latent embeddings of all the labeled data with respect to the intermediate model, namely $\{\bm{z}=h_{\hat{\theta}}(\bm{x}):\bm{x}\in\mathcal{S}_{l}\}$, and then apply k-means clustering algorithm to partition the embeddings into $k$ clusters. Note that this step is different from LCS-KM, because we now do not pre-generate the unlabeled dataset. Let $\{\bm{\mu}_{1},\bm{\mu}_{2},\ldots,\bm{\mu}_{k}\}$ denote the centroids corresponding to the $k$ returned clusters. We define the guidance loss as:
[…]
where $d_{1}(\bm{x};h_{\hat{\theta}})=\min_{j\in[k]}\|h_{\hat{\theta}}(\bm{x})-\bm{\mu}_{j}\|_{2}$ denotes the Euclidean distance between the latent embeddings of $\bm{x}$ and the closest centroid, $d_{2}(\bm{x};h_{\hat{\theta}})=\min_{j\in[k],j\neq j_{1}}\|h_{\hat{\theta}}(\bm{x})-\bm{\mu}_{j}\|_{2}$ is the second smallest distance, and $j_{1}$ stands for the index of the closest centroid. By definition, we know $\ell_{\mathrm{KM}}(\bm{x})$ is always non-negative, and a smaller value of $\ell_{\mathrm{KM}}(\bm{x})$ indicates that the input $\bm{x}$ is more likely to be a boundary-adjacent point.
LCG-GMM. Similarly, we can define an informative guidance loss based on the coefficients of the Gaussian mixture models. Let $\{\bm{\mu}_{j},\mathbf{\Sigma}_{j})$ be the mean and covariance parameters returned by GMM with respect to the latent embeddings $\{\bm{z}=h_{\hat{\theta}}(\bm{x}):\bm{x}\in\mathcal{S}_{l}\}$. Let $p_{j}(\bm{z})$ be the $j$-th GMM posterior probability defined by Equation 7. Then, the guidance loss is defined as:
[…]
Guided DDPM Fine-Tuning. To utilize our designed guidance losses in DDPM fine-tuning, we propose to first denoise $\bm{x}_{t}$ via a single DDPM reverse step $g_{\theta}(\bm{x}_{t},t)$, where $\bm{x}_{t}$ is a noisy image at $t$-th time step defined by the forward process, and then apply the respective guidance loss to $g_{\theta}(\bm{x}_{t},t)$. For any sample $\bm{x}_{0}\sim q_{0}$, time $t\sim\mathcal{U}(T)$ and noise $\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$, we define the following loss function $L_{\mathrm{reg}}(\theta)$ that will be used as an additional regularization term to fine-tune DDPM:
$L_{\mathrm{reg}}(\theta)=\mathbb{E}_{\bm{x}_{0},t,\bm{\epsilon}}\big[\ell_{\mathrm{guide}}\big(g_{\theta}(\sqrt{\bar{\alpha}_{t}}\bm{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon},t)\big)\big],$ (15)
where $\ell_{\mathrm{guide}}$ corresponds to Equation 12 when we use prediction confidence, and refers to Equations 13 and 14 when using LCG-KM and LCG-GMM, respectively. Putting all the pieces together, the fine-tuning loss for direct generation of boundary-adjacent data is defined as:
$L_{\mathrm{tot}}(\theta)=L_{\mathrm{DDPM}}(\theta)+\lambda\cdot L_{\mathrm{reg}}(\theta),$ (16)
where $\lambda\geq 0$ balances the trade-off between high-quality image generation captured by the original DDPM loss and boundary-seeking data generation driven by our guidance loss. By penalizing the model weights through the regularization term, our method gradually pushes $\theta$ towards higher probabilities of generating boundary-adjacent data using stochastic gradient descent (SGD). Our later experiments show that when starting from a pre-trained DDPM (i.e., on CIFAR-10), a few fine-tuning epochs are sufficient to bias the diffusion model to generate the desirable subset of synthetic images.
IV Experiments
In this section, we test the performance of our data reduction strategies proposed in Section III compared with baseline SSAT algorithms. In particular, we first conduct experiments to evaluate our selection-based methods in terms of robustness and efficiency advantages when integrated into the SSAT pipeline (Section IV-A). Then, we highlight the computational advantages of our approaches, particularly their capabilities of accelerating the convergence and reducing the total runtime of SSAT (Section IV-B). Finally, we study the effectiveness of our DDPM fine-tuning schemes in further improving the efficiency of SSAT algorithms under a generative setup (Section IV-C). To begin with, we lay out the main experimental setup. All our experiments are conducted using a single A100 GPU. All the remaining details, including all the hyperparameters, for reproducing our experiments are provided in Appendix B.
Dataset. We test two image benchmarks in our evaluation, SVHN [26] and CIFAR-10 [27], both widely used in prior adversarial ML literature. We consider two scenarios in terms of the source of the utilized unlabeled dataset $\mathcal{S}_{u}$ for testing our selection-based strategies: (i) external–using an external dataset, and (ii) generative–using a synthetic dataset generated by some pre-trained generative model. For external scenarios, we follow the implementation protocols outlined by Carmon et al. [11]. To be more specific, SVHN models are trained on 7373K labeled digit images from the original dataset, supplemented by 531531K extra unlabeled SVHN images. For CIFAR-10, the original dataset comprises 5050K labeled training images and 1010K labeled test images, where models trained using SSAT are augmented by an external unlabeled dataset $\mathcal{S}_{u}$ with 500500K images sampled from the 8080M Tiny Images (80M-TI) dataset. For generative setup, we adopt 11M synthetic images generated by DDPM for both datasets, following the protocols documented by Gowal et al. [13].
Configuration. Following prior literature, we mainly consider $L_{\infty}$-norm bounded perturbations with $\epsilon=0.015$ on SVHN and $\epsilon=0.031$ on CIFAR-10. We adopt WideResNet (WRN) architectures and train models using TRADES [9] on both labeled and unlabeled data, coupled with pseudo labels returned by the intermediate model. In Appendix D, we evaluate the generalizability of our methods for varying $\ell_{\infty}$ perturbation size, vanilla PGD-based adversarial training with varying attack steps, and $\ell_{2}$ perturbations. To standardize training, an epoch is defined as processing 5050K data points, regardless of the total dataset size. Prior works also adopted this approach [11, 13], which ensures that training time per epoch remains consistent across datasets of varying sizes.
SSAT. For experiments with external unlabeled data, SSAT algorithms are trained for a total of $200$ epochs, whereas experiments with generated data extend to $400$ epochs. To compare the performance of our methods in a fair way, models are saved every $25$ epoch of SSAT, and the model achieving the highest robust accuracy is selected as the best model. This approach aligns with the early stopping practices commonly used in adversarial ML literature [9, 28]. In experiments with external data and ratio parameter $\alpha\in\{10\%,20\%\}$, we consistently observe peak robust accuracy around $100$ epochs for CIFAR-10 and $75$ epochs for SVHN. Conversely, when no data selection scheme is applied, peak performance is delayed to approximately $200$ epochs. Total training time is reported as the duration required to achieve the best model performance.
Intermediate Model. Our data selection and guided diffusion schemes require training an intermediate model $f_{\hat{\theta}}$, which we adopt the same architecture as the final model. This intermediate model is trained using standard supervised learning for $100$ epochs on a WRN-28-10 architecture, which takes around $55$ minutes. Notably, this time is excluded from the reported total runtime. The intermediate model fulfills two critical roles: it not only facilitates the pseudo-labeling of unlabeled data but also constitutes a necessary component for implementing any of our proposed strategies. As such, this training step is integral to the overall workflow, irrespective of whether a specific selection or fine-tuning method is employed.
Evaluation Metric. To test the robustness of the final model returned by SSAT, we mainly use multi-step PGD attacks (PGD) [7], due to their broad adoption in adversarial ML. We also report robust accuracies against AutoAttack (AA) [29] for more rigorous robustness evaluation and the model’s clean accuracy on normal inputs (Clean) as a reference.
TABLE I: SSAT performance under various unlabeled data selection schemes. We consider $\ell_{\infty}$ perturbations with $\epsilon=0.015$ for SVHN, and $\ell_{\infty}$ perturbations with $\epsilon=0.031$ for CIFAR-10. For each configuration, we underline the best robust accuracies.
Dataset Ratio $\bm{\alpha}$ Method External Unlabeled DDPM-Generated
Clean ($\%)$ PGD ($\%)$ AA ($\%)$ Time (h) Clean ($\%)$ PGD ($\%)$ AA ($\%)$ Time (h)
SVHN $0\%$ No Sel. $94.4$ $75.6$ $68.3$ $2.62$ $94.4$ $75.6$ $68.3$ $2.62$
$100\%$ No Sel. $97.1$ $86.0$ $75.3$ $8.38$ $97.4$ $86.3$ $75.2$ $17.41$
$1\%$ Random $94.2$ $78.4$ $70.3$ $3.28$ $95.1$ $80.3$ $70.5$ $3.28$
PCS $95.4$ $81.8$ $71.6$ $3.75$ $95.4$ $81.5$ $71.5$ $4.08$
LCS-GMM $95.1$ $81.8$ $72.1$ $3.97$ $95.6$ $81.2$ $71.8$ $4.27$
LCS-KM $95.2$ $\underline{82.7}$ $\underline{72.9}$ $3.90$ $96.4$ $\underline{82.4}$ $\underline{72.6}$ $4.18$
$10\%$ Random $95.3$ $82.0$ $73.0$ $3.28$ $95.6$ $83.6$ $72.4$ $3.28$
PCS $96.1$ $82.8$ $74.2$ $3.85$ $95.3$ $84.0$ $73.5$ $4.08$
LCS-GMM $96.2$ $83.0$ $74.3$ $3.97$ $95.5$ $84.1$ $74.1$ $4.27$
LCS-KM $96.1$ $\underline{86.3}$ $\underline{75.2}$ $3.90$ $96.6$ $\underline{86.6}$ $\underline{74.8}$ $4.18$
$20\%$ Random $96.2$ $82.3$ $73.2$ $3.28$ $96.3$ $84.6$ $74.0$ $3.28$
PCS $96.3$ $83.9$ $74.7$ $3.75$ $96.6$ $85.4$ $74.5$ $4.08$
LCS-GMM $96.4$ $84.3$ $75.0$ $3.97$ $95.8$ $85.5$ $74.7$ $4.27$
LCS-KM $96.2$ $\underline{86.6}$ $\underline{75.1}$ $3.90$ $96.7$ $\underline{87.2}$ $\underline{75.3}$ $4.18$
CIFAR-10 $0\%$ No Sel. $84.9$ $55.4$ $49.2$ $10.69$ $84.9$ $55.4$ $49.2$ $10.69$
$100\%$ No Sel. $89.7$ $62.5$ $58.6$ $28.50$ $85.7$ $61.8$ $58.4$ $57.13$
$1\%$ Random $83.2$ $54.5$ $49.8$ $14.25$ $83.3$ $54.3$ $49.6$ $14.25$
PCS $84.9$ $55.4$ $51.5$ $14.73$ $84.3$ $55.2$ $50.7$ $15.16$
LCS-GMM $84.2$ $55.1$ $52.1$ $14.89$ $85.8$ $54.6$ $51.8$ $15.36$
LCS-KM $85.6$ $\underline{56.4}$ $\underline{52.9}$ $14.78$ $85.9$ $\underline{56.1}$ $\underline{52.5}$ $15.27$
$10\%$ Random $85.2$ $56.0$ $52.4$ $14.25$ $85.0$ $56.5$ $52.6$ $14.25$
PCS $85.3$ $56.9$ $53.7$ $14.73$ $85.4$ $56.9$ $54.1$ $15.16$
LCS-GMM $85.9$ $57.1$ $54.2$ $14.89$ $85.8$ $57.1$ $54.5$ $15.36$
LCS-KM $87.2$ $\underline{58.2}$ $\underline{55.3}$ $14.78$ $86.1$ $\underline{58.0}$ $\underline{55.8}$ $15.27$
$20\%$ Random $87.1$ $57.5$ $54.2$ $14.25$ $85.4$ $57.2$ $54.2$ $14.25$
PCS $87.0$ $57.9$ $54.5$ $14.73$ $85.6$ $58.0$ $55.0$ $15.16$
LCS-GMM $87.0$ $58.2$ $55.5$ $14.89$ $85.9$ $58.6$ $55.8$ $15.36$
LCS-KM $88.7$ $\underline{60.7}$ $\underline{57.8}$ $14.78$ $85.5$ $\underline{60.2}$ $\underline{57.2}$ $15.27$
IV-A Strategic Unlabeled Data Selection
SVHN. Our evaluation results on SVHN are summarized in Table I, which demonstrate the efficacy of the proposed data selection schemes when employed in SSAT algorithms trained using a WRN-16-8 architecture with varying data ratios $\alpha\in\{1\%,10\%,20\%\}$, with comparisons to baseline SSAT performance with $\alpha\in\{0\%,100\%\}$. Specifically, employing our LCS-KM strategy with $10\%$ or $20\%$ of extra unlabeled data achieves robust accuracies of $86.3\%$ and $86.6\%$, respectively, which matches and even slightly outperforms the $86.0\%$ robustness achieved using the entire unlabeled dataset. In contrast, random selection with $20\%$ of the data yields a robust accuracy of only $82.3\%$, underscoring the superior efficacy of informed selection strategies such as LCS-KM. The robust accuracies evaluated using AutoAttack (i.e., the AA column of Table I) demonstrate similar patterns compared to those using multi-step PGD Attacks (i.e., the PGD column of Table I). Additionally, alternative methods, including PCS and LCS-GMM, consistently outperform random selection, confirming the effectiveness of our insights. Notably, the robustness of models trained using SSAT peaks when the boundary ratio parameter $\beta$ is inversely proportional to the selection ratio $\alpha$, suggesting that smaller $\beta$ values are favorable as the data inclusion proportion increases. We provide detailed sensitivity analyses of these hyperparameters in Section V-C.
Our evaluations with 11M DDPM-generated data highlight that when using $10\%$ to $20\%$ of unlabeled data selected via LCS-KM, SSAT can achieve comparable robustness to the SSAT baseline. Since all SSAT algorithms listed under the column “DDPM-Generated” in Table I require the same amount of pre-generation time, we exclude it from the total runtime. In particular, training SSAT with reduced data requires only $75$ epochs to achieve the peak robust accuracy, while training with the full dataset requires up to $400$ epochs. Such a large reduction in computation time underscores the practical advantages of the proposed data selection strategies, which are further elaborated in Section IV-B. Besides, we observe in our experiments that using ground-truth labels reveals negligible differences compared to pseudo-labeled data, indicating that the primary performance gains stem from the selection strategy itself rather than label accuracy (see Table IV in Appendix C).
CIFAR-10. We adopt a WRN-28-10 architecture for CIFAR-10 experiments. To assess the generalizability of our findings, we also conduct ablations on ResNet-18, which are detailed in Table V in Appendix C. Across varying ratios of utilized unlabeled data $\alpha$ and different data selection schemes, Table I reveals that selecting $20\%$ of the external data using LCS-KM attains robust accuracy of $60.7\%$, comparable to the $62.5\%$ obtained with the entire dataset. In comparison, random selection with the same ratio parameter achieves only $57.8\%$, underscoring the advantages of our strategic LCS technique with k-means. Similar to SVHN experiments, we extend our analysis to CIFAR-10 with synthetically generated data. Using 11M DDPM-generated CIFAR-10-like images, we discover that using just $20\%$ of the generated data via LCS-KM results in performance nearly identical to that achieved with the full dataset. These trends are also consistent across experiments with $\ell_{2}$ perturbations (see Figure 10 for supporting evidence).
Our findings emphasize the robustness and adaptability of the proposed data selection schemes, especially for LCS-KM, even under varying experimental configurations. In terms of computational efficiency, CIFAR-10 experiments highlight the significant advantages of data selection. When leveraging selected external unlabeled data, optimal robust performance was achieved after $100$ epochs of training. By contrast, using the full unlabeled or generated datasets without selection required extended training durations of $200$ and $400$ epochs, respectively, to reach comparable performance. These results validate the efficacy of our latent clustering-based selection approach in enhancing semi-supervised adversarial training. The proposed method not only maintains the robust accuracy of the final model, $\theta_{\mathrm{final}}$, but also substantially reduces computational costs and accelerates the training process, underscoring its practical utility and scalability in SSAT frameworks.
(a) Vanilla AT
(b) Random Selection ($\alpha=10\%$)
(c) LCS-KM ($\alpha=10\%$)
(d) Full Utilization ($\alpha=100\%$)
Figure 1: Illustration of standard and robust accuracy curves of SSAT on CIFAR-10 with different configurations of unlabeled data selection from the external 500500K Tiny Images: (a) No extra data, (b) random selection with $\alpha=10\%$, (c) LCS-KM with $\alpha=10\%$, and (d) utilizing all 500500K unlabeled data.
TABLE II: Runtime comparison of SSAT with synthetically generated unlabeled data across various settings on CIFAR-10.
Ratio $\bm{\alpha}$ Method Fine-Tuning (h) Generation (h) Selection (h) #Epochs Total Time (h) PGD (%)
$0\%$ No Gen. $100$ $10.7$ $55.4$
$100\%$ Std. DDPM $3.90$ $400$ $61.0$ $61.8$
$20\%$ LCS-KM $3.90$ $0.97$ $100$ $19.1$ $60.3$
$20\%$ LCG-KM $0.43$ $0.77$ $100$ $15.7$ $60.2$
TABLE III: Performance of our DDPM fine-tuning strategies on CIFAR-10 under varying $\alpha$. The best robust accuracies are underlined.
Ratio $\bm{\alpha}$ Method Clean ($\%$) PGD ($\%$) AA ($\%$)
$0\%$ No Gen. $84.9$ $55.4$ $49.2$
$100\%$ Std. DDPM $85.7$ $61.8$ $58.4$
$1\%$ PCG $84.4$ $55.4$ $52.8$
LCG-GMM $85.3$ $55.8$ $52.6$
LCG-KM $85.2$ $56.5$ $53.1$
$10\%$ PCG $84.7$ $56.7$ $54.2$
LCG-GMM $85.8$ $57.6$ $55.0$
LCG-KM $84.2$ $57.9$ $55.6$
$20\%$ PCG $85.9$ $58.3$ $55.9$
LCG-GMM $84.7$ $58.3$ $56.1$
LCG-KM $85.7$ $60.3$ $57.4$
IV-B Computational Benefits of Data Reduction
As suggested by the runtime comparisons demonstrated in Table I, utilizing smaller, carefully selected unlabeled data can significantly reduce the overall training time of SSAT while maintaining competitive robustness. We further visualize the learning curves of SSAT and discuss the computational benefits associated with our proposed methods.
Figure 1 depicts the learning curves of vanilla AT and SSAT algorithms on CIFAR-10 with varying ratios of unlabeled data selected from TinyImages. In our setup, one training epoch corresponds to processing 5050K data points, with $10$ PGD steps for training and $20$ for testing. When no additional data is utilized, the highest robust test accuracy is achieved after approximately $75$ epochs (Figure 1a). Incorporating $10\%$ additional data, whether through random selection or curated approaches, extends the optimal convergence point to approximately $100$ epochs (Figures 1b and 1c). In contrast, leveraging $100\%$ of the unlabeled data delays the achievement of peak accuracy to around $185$ epochs (Figure 1d). These findings suggest that including extra unlabeled data in SSAT increases computational demands. Our selection schemes reduce redundant information in the dataset, allowing SSAT to achieve high robustness performance with fewer training epochs.
Generally speaking, the size of the utilized unlabeled dataset is critical in determining the time required for SSAT to achieve peak robust accuracy. Larger unlabeled datasets usually extend the convergence times, whereas smaller datasets facilitate faster convergence, allowing the model to reach optimal accuracy earlier in training. However, continued training beyond this point often leads to robust overfitting. Early stopping offers a computational advantage in scenarios with limited datasets by preventing overfitting with reduced training durations. Nevertheless, early stopping proves less effective in SSAT when working with large-scale unlabeled datasets, as longer training is necessary to achieve peak performance. To mitigate these challenges, selecting a smaller yet strategically curated subset of data significantly reduces the dataset size, enabling faster convergence and minimizing overall training time. These findings underscore the computational benefits of our proposed data selection schemes, providing a practical solution to the challenges of SSAT with extensive datasets.
(a) Random Selection ($\alpha=10\%$)
(b) LCS-KM ($\alpha=10\%$)
(c) LCS-KM ($\alpha=20\%$)
(d) Full Utilization ($\alpha=100\%$)
Figure 2: Standard and robust accuracy curves of SSAT with labeled data from COVIDGRand unlabeled data from CoronaHack, with different selection schemes and ratios: (a) random selection with $\alpha=10\%$, (b) LCS-KM with $\alpha=10\%$, (c) LCS-KM with $\alpha=20\%$, and (d) all unlabeled data $\alpha=100\%$.
IV-C Guided DDPM Fine-Tuning
We study the performance of our guided DDPM fine-tuning methods as depicted in Algorithm 3 on CIFAR-10. Unlike our previous selection schemes, which required pre-generating the whole dataset, here we generate the same amount of unlabeled data matching the ratio $\alpha$. The data-mixing hyperparameters, namely $\alpha$ and $\beta$, are set consistently with those used in strategic selection. We employ a pretrained DDPM model trained on CIFAR-10 for these guided fine-tuning experiments. Table III presents the results across different $\alpha$ ratios, where our DDPM fine-tuning schemes achieve competitive clean accuracy and robustness performance, matching their selection counterpart. For instance, under the ratio scenario $\alpha=20\%$, leveraging LCG-KM to generate boundary-adjacent data achieves a PGD robust accuracy of $60.3\%$ when integrated in the SSAT framework, which matches the result of $60.2\%$ obtained by LCS-KM. Again, latent clustering with k-means consistently achieves the best robustness compared to the other alternatives. These findings mirror our findings from strategic selection, reinforcing that a small but informative generated set can yield performance comparable to a much larger dataset.
Moreover, we conduct a detailed runtime analysis across different data reduction schemes under the generative setup. In particular, we choose to present the best-performing methods with the optimal hyperparameters, namely LCS-KM and LCG-KM under the scenario of $\alpha=20\%$, and compare the end-to-end training times. We also report the baseline performance of SSAT, where the whole generated dataset ($\alpha=100\%$) and no generated data ($\alpha=0\%$) are used, respectively. The comparison results are summarized in Table II. For clarity, we separately report the computational time of DDPM fine-tuning, data generation, data selection, and the total SSAT training process. Our DDPM generator’s throughput is roughly $71$ images per second, so the approximate data generation times are 200200K images in $46$ minutes, 11M images in $3.9$ hours, and 100100M images in $16.2$ days. As a result, our guided DDPM finetuning (LCG-KM) approach reduces the total runtime from $19.1$ hours to $15.7$ hours when compared to LCS-KM. The improvement will be more pronounced when the total amount of data generated by DDPM is much larger. These results highlight the trade-off between robustness gains and computational cost. While generating millions of synthetic samples can enhance robustness, it is computationally expensive. Our guided fine-tuning approach offers a more resource-efficient alternative, reducing SSAT total runtime from $61.0$ hours to $15.7$ hours while preserving strong robustness.
(a) PCS
(b) LCS-GMM
(c) LCS-KM
Figure 3: Visual comparison of selection techniques on TinyImages dataset in the latent space. Each subplot represents a different method: (a) PCS identifies points with the lowest classification confidence, highlighting areas where the model is most uncertain within the ten-class latent representation, (b) LCS-GMM illustrates probability contours from Gaussian Mixture Models, with selected points emphasizing regions of overlapping probabilities among the ten class clusters, and (c) LCS-KM highlights points selected near decision boundaries across ten classes based on k-means clustering in the latent space.
(a) CIFAR-10
(b) SVHN
(c) CIFAR-10
(d) SVHN
Figure 4: Illustration of SSAT performance using our LCS-KM method on image benchmarks with varying boundary and unlabeled data reduction ratio parameters with respect to: (a) and (c) DDPM-generated CIFAR-10 data with $\alpha=20\%$, (b) and (d) 531531K extra SVHN data with $\alpha=10\%$. For each figure, we vary the considered ratio parameter while keeping all the remaining hyperparameters fixed.
V Additional Analyses
In this section, we present additional analyses to provide further insights into our methods. We first evaluate their performance on a real-world medical application dataset (Section V-A), and then visualize the selected data distribution in the latent space to qualitatively understand their ability in identifying critical data points (Section V-B). Finally, we provide ablation studies on key hyperparameters of our methods. Due to page limit, additional experiments on other hyperparameters, generalizability under different training schemes and perturbations, and the effect of intermediate model quality in low-label regimes, are provided in Appendix D.
V-A Application to Real-World Medical Data
To test the generalization of our findings, we extend our study to a real-world healthcare application. COVIDGR [30] is used as the labeled dataset, and CoronaHack [31] serves as the source of the unlabeled data. Specifically, COVIDGR contains $852$ chest X-ray images, evenly distributed across two classes: $426$ COVID-19 positive (RT-PCR confirmed) and $426$ negative cases. It is then split into $80\%$ training and $20\%$ test sets. CoronaHack comprises $5910$ images with an imbalanced class distribution ($73\%$ positive, $27\%$ negative), split into an $89\%$ training set and an $11\%$ test set. We consider $\ell_{\infty}$ perturbations with $\epsilon=0.1$ and train models on a ResNet-18 architecture using SSAT by applying our LCS-KM strategy to select unlabeled CoronaHack data with varying ratios. Further details of the experimental setup is discussed in Appendix B.
The results are demonstrated in Figure 2, which confirms the effectiveness of our selection scheme in enhancing robustness while optimizing efficiency. For instance, using LCS-KM to select $10\%$ of the unlabeled data achieves a robust accuracy of $56\%$, compared to $53\%$ when the same proportion is chosen randomly (Figures 2a and 2b). In comparison, using a smaller unlabeled dataset leads to faster convergence, underscoring the benefits of early stopping. Specifically, the best performance of $56\%$ is reached at epoch $36$ when using $10\%$ of the data selected with LCS-KM (Figure 2b). Figure 2c shows that incorporating $20\%$ of the data with LCS-KM improved the accuracy to $57\%$ at epoch $47$. In contrast, training on the entire unlabeled dataset requires $60$ epochs to achieve the peak robust accuracy of $58\%$ (Figure 2d). These findings confirm that using a smaller but strategically selected unlabeled data subset can achieve robustness comparable to that of the whole dataset while significantly reducing computational costs. These results highlight the practical relevance of our data reduction scheme, particularly in medical diagnostics, where labeled data is limited, yet robust model performance is essential.
V-B Comparative Visualizations of Our Methods
To facilitate a deeper understanding, we perform a comparative visualization of the selected unlabeled data on CIFAR-10. We first extract their latent representations from the penultimate layer of a pre-trained WRN-28-10 intermediate model. For visualization, we employ a two-step dimension reduction that balances the retention of global structure and the visualization of local relationships. Specifically, we first perform Principal Component Analysis (PCA) [32]. By reducing the dimensionality of the latent features to $50$, PCA retains the majority of the variance within the data. In the second step, we apply t-Distributed Stochastic Neighbor Embedding (t-SNE) [33] to project the $50$-dimensional PCA-reduced features onto a two-dimensional space. Unlike PCA, which emphasizes preserving variance, t-SNE mainly aims to maintain the local relationships among data points. After projecting onto the two-dimensional space, we visualize the set of unlabeled data returned by our selection schemes under $\alpha=10\%$.
Figure 3 presents the visualization results, where each color represents the label class predicted by the intermediate model for the corresponding unlabeled point, and the highlighted black dots indicate the data points selected by the respective selection strategy. Our visualization results largely explain why our LCS-KM method consistently achieves the best robustness performance when incorporated in SSAT. In particular, PCS adopts a purely confidence-based selection strategy, choosing the globally lowest confidence points without considering their distribution within or across classes. Although this direct approach is appealing for its simplicity, it can inadvertently prioritize noisy outliers that offer limited benefit for refining the decision boundary. As the number of classes grows and data complexity increases, this scattershot selection tends to yield a more disorganized set of points (Figure 3a).
In comparison, LCS-GMM and LCS-KM yield more structured sampling patterns. While LCS-GMM can roughly capture regions where classes overlap, it assumes that each cluster follows approximately Gaussian distributions, an assumption that may fail in real-world settings. As a result, LCS-GMM’s contours become misaligned with the true underlying structure, leading to less accurate localization of essential boundary points (Figure 3b). A closer look at the selected samples also reveals that the corresponding region is broader and less concentrated on the decision boundary, making it harder to capture truly critical examples for training. In some instances—such as the cyan region in the figure—LCS-GMM fails to cover as many relevant boundary samples as LCS-KM.
Finally, our best-performing LCS-KM method emphasizes the importance of leveraging the latent-space clustering to identify boundary-adjacent data points in a more accurate and balanced manner. By partitioning the latent embeddings into multiple k-means clusters, LCS-KM is able to locate more points with low classification confidence within each cluster (Figure 3c). Compared with LCS-GMM, a notable pattern is that the unlabeled data picked by LCS-KM are aligned more linearly. This coincides with the piecewise linear characteristic of neural networks’ decision boundary, especially for the last layer. We hypothesize that such a local linear property potentially explains why our latent clustering-based approach with k-means always achieves the best performance compared to the other two alternatives, which is worth further investigation.
V-C Hyperparameter Sensitivity
The success of our methods relies on a few key hyperparameters, including $\beta$ that controls the balance between boundary-adjacent unlabeled data and the remaining data points, the proportion of selected unlabeled data $\alpha$, and the per-batch ratio of extra to original data $\gamma$. We record model robust accuracy by varying each parameter to provide guidelines on how to choose these hyperparameters optimally. Due to page limits, the detailed settings of these experiments and the ablations on $\gamma$ are deferred to Appendix C-A.
Ratio of Boundary Data $\bm{\beta}$. Concentrating exclusively on data points near the decision boundary can result in an excessive number of points close to the boundary and too few farther away, leading to overfitting. To mitigate this, we choose a combination of points, with some near the boundary and others further away, where the ratio is determined by $\beta$. Figure 4a illustrates how robust accuracy varies with $\beta$ on CIFAR-10 using $20\%$ generated data from DDPM. In this scenario, the optimal robust accuracy is achieved when $\beta$ is $0.4$, meaning $40\%$ of the selected points are near the boundary, while $60\%$ are farther away. Figure 4b illustrates how robust accuracy varies with $\beta$ on SVHN using $10\%$ extra data from the same dataset. Due to a smaller amount of external data, selecting points near the boundary results in a lower risk of overfitting. As a result, the highest robust accuracy is observed when $\beta$ is $0.6$ with $60\%$ of the points near the boundary and $40\%$ farther away. Therefore, the ideal value of $\beta$ depends on the amount of external data added and the ratio of external to original data. When more external data is added, using a smaller $\beta$ usually leads to better robust accuracy. Conversely, choosing a larger $\beta$ is favorable with less extra unlabeled data.
Data Reduction Ratio $\bm{\alpha}$. Figures 4c and 4d describe how accuracies vary with the ratio parameter $\alpha$. We analyze how the quantity of extra unlabeled data influences the effectiveness of our data reduction method, which is designed to prioritize points near the decision boundary. When the available data is limited, our algorithm excels by selecting a higher proportion of boundary-adjacent data, achieving a balanced selection that enhances robust accuracy. However, as the volume of unlabeled data increases, the algorithm’s focus on boundary points can inadvertently lead to overfitting, diminishing the improvement gained on the model’s generalization ability. Interestingly, as the data volume continues to grow and more points are added, the performance begins to improve once again, suggesting that the initial overfitting is mitigated by the sheer abundance of data, leading to a recovery in accuracy.
(a) Fine-Tuning Epochs $S$
(b) Regularization Strength $\lambda$
Figure 5: Clean and robust accuracy curves of SSAT with LCG-KM using generated data on CIFAR-10 under $\alpha=10\%$ by varying: (a) the number of fine-tuning epochs $S$, and (b) the regularization strength parameter $\lambda$.
Guided DDPM Finetuning. We conduct ablation studies on our guided DDPM fine-tuning approach to better illustrate its behavior. In particular, we focus on two key hyperparameters, the number of fine-tuning epochs $S$ and the regularization strength parameter $\lambda$, and study their impact on accuracy and robustness performance when training SSAT algorithms on CIFAR-10 with LCG-KM and $\alpha=10\%$. We also investigate the quality of the data samples generated by our fine-tuned DDPM using both quantitative and qualitative metrics.
Figure 5a illustrates our ablation results on $S$. We vary the number of fine-tuning epochs $S\in\{5,15,25,35,45\}$ with $\lambda=0.5$, and report the final clean and robust accuracies after the whole SSAT process. Figure 6 visualizes the images generated by our DDPM models with different fine-tuning epochs. We observe that choosing a small number of epochs for fine-tuning DDPM (e.g., 10102020 epochs) is usually sufficient to achieve the best robustness performance, likely because the starting DDPM is already pre-trained for CIFAR-10. Training for too many epochs can lead to a degradation in image quality and loss of diversity, thereby adversely influencing the final robust accuracies. In addition, Figure 5b shows the results for the regularization strength parameter $\lambda\in\{0.1,0.3,0.5,0.7\}$ with fixed epoch number $S=15$. Our findings show that choosing $\lambda$ within the range of $[0.4,0.6]$ results in the best performance, striking a balance between the quality of the generated images and their proximity to the model’s decision boundary.
Figure 6: Visualization of generated data across different fine-tuning epochs
VI Conclusion and Future Work
In this work, we proposed various unlabeled data reduction schemes to improve the efficiency of SSAT while preserving robust accuracy. In particular, we emphasized that strategically selecting or generating more boundary-adjacent data points can significantly improve the efficiency of semi-supervised adversarial training algorithms, without losing their robustness advantages. Among our proposed methods, we found that latent-based approaches using k-means clustering, including LCS-KM and LCG-KM, perform the best, as they reduce computational costs and memory requirements.
Limitation & Future Work. We acknowledge that our work has several limitations. While the selection scheme with LCS-KM consistently achieves the best performance, its success requires carefully choosing a few hyperparameters (e.g., to avoid overfitting). Future studies can leverage AutoML algorithms to automatically determine these hyperparameters. Additionally, it remains unclear whether our best-performing latent-clustering-based methods (i.e., LCS-KM and LCG-KM) can be further optimized to achieve better performance. Theoretical analysis of how to solve the optimization problems (Equations 4 and 6) in a more principled way would be helpful. Furthermore, our guided DDPM fine-tuning approach proposes to adapt a pre-trained model for boundary-seeking data generation. An interesting future direction is to investigate whether training-free sampling-based guidance [34] or more advanced generative models [35, 36] can enhance the efficiency and robustness of SSAT. Despite these limitations, we believe our work can serve as an important initial step towards the more efficient and scalable development of robust models.
Availability
The implementation of the latent-based data reduction methods and all our experiments are available as open-source code at: https://github.com/TrustMLRG/EfficientSSAT.
Appendix A Related Work
A-A Adversarially Robsut Learning
Adversarial examples are inputs crafted with small perturbations that are designed to mislead model predictions [1]. The prevalence of adversarial examples with deep neural networks poses a critical challenge in modern machine learning, especially for security-critical or safety-critical applications. Earlier-proposed attacks like fast gradient sign method (FGSM) [2] and projected gradient descent (PGD) [7] introduced gradient-based techniques to find adversarial examples using constrained optimization, while Carlini and Wagner proposed unconstrained CW attacks [37], which search for the minimal perturbation that causes misclassification. Recent attacks like Auto-PGD [38] automate gradient steps for higher efficiency in finding worst-case perturbations, while AutoAttack [17] is an ensemble of diverse parameter-free attacks, making them more effective in evaluating model robustness.
Adversarial Training. On the defense side, adversarial training and its variants are the most popular methods for achieving robustness. For instance, Madry et al. proposed to train DNNs on adversarially perturbed inputs using PGD attacks [7], while Zhang et al. introduced TRADES [9], which balances the trade-off between robustness and standard accuracy. Nevertheless, adversarial training algorithms are often criticized for overfitting adversarially perturbed samples produced during training time and for their high computation costs, which hinder their deployment in real-world applications. In particular, Rice et al. illustrated through comprehensive experiments that the robust overfitting phenomenon is prevalent in adversarial training [28], where early stopping is shown to be effective in alleviating such an issue. Techniques, such as Free AT [18] and Fast AT [19], offer promising solutions to reduce the computation for adversarial training. Nevertheless, they suffer from decreased robustness performance.
Moreover, a line of research pointed out that treating all the data points equally in adversarial training is not optimal for robust learning since different examples can impact model robustness differently. For example, Zhang et al. argued that data points far from decision boundaries are inherently more secure and less likely to be influential in robust learning against adversarial perturbations [16]. As a result, these points should be given less weight during training to prevent the model from overfitting them, which could negatively affect its generalization on more vulnerable or unseen data. In a similar vein, Hua et al. advocated for a more targeted approach [24], recommending that PGD training be applied primarily to examples situated near the decision boundaries. This strategy focuses the model’s attention on the points that are most likely to enhance its robustness, as adversarial perturbations on boundary-adjacent data are more likely to result in meaningful improvements to the model’s resistance to attacks. Our work builds on these ideas by strategically identifying and prioritizing boundary-adjacent points from the unlabeled dataset to address the unique challenges associated with self-supervised adversarial training methods.
Self-Supervised Adversarial Training. Utilizing unlabeled data to improve model performance has become an active and rapidly advancing field of research [39]. In the adversarial context, Schmidt et al. [10] highlighted a key observation: the sample complexity required for robust learning far exceeds that of standard learning. They showed that achieving adversarially robust generalization often necessitates a much larger dataset than traditional learning approaches, posing a challenge when labeled data is scarce or costly. Subsequently, a line of works proposed to boost model robustness by involving additional unlabeled data, often acquired from a similar but slightly different distribution, using self-supervised methods in adversarial training [11, 40, 41, 12]. In particular, Carmon et al. proposed robust self-training methods to address the sample complexity issue [11], which first trains an intermediate model using available labeled data and then leverages the model to generate pseudo-labels for unlabeled data. The pseudo-labeled samples, combined with the original labeled data, are subsequently used to train a final robust model. Concurrently, Najafi et al. proposed an optimization framework incorporating labeled and unlabeled data [41], while providing formal guarantees on adversarial robustness.
Leveraging synthetic data produced by state-of-the-art generative models to build robust models has also been explored extensively in recent literature [13, 14, 15]. For instance, Gowal et al. employed various unconditional generative models [13], including variational autoencoders (VAEs), generative adversarial networks (GANs), and the more advanced denoising diffusion probabilistic models (DDPMs), to produce a large synthetic dataset to be incorporated in adversarial training, which improves robustness further. While all the aforementioned SSAT methods enrich the training dataset, they require considerable additional unlabeled data to ensure effective robustness enhancement. In this work, we explore data selection schemes to reduce the amount of unlabeled or generated data that are both efficient and effective for self-supervised adversarial training to achieve good robustness performance.
A-B Data-Efficient Deep Learning
Within the broader literature of data-efficient deep learning, various techniques have been proposed to reduce the amount of training data while retaining critical information. Two prominent approaches are dataset distillation and coreset selection.
Dataset Distillation. Dataset distillation aims to distill large datasets into compact, synthetic, yet informative datasets by matching gradients or optimizing specific loss functions, allowing models to achieve comparable performance with significantly fewer data points. For instance, Wang et al. proposed a foundational approach to dataset distillation [42], proposing methods to condense large datasets into smaller, synthetic subsets while preserving model performance. Later, Zhao et al. introduced the concept of gradient matching [43], where distilled data points are optimized to match the gradients of the original dataset. This technique largely improves the distillation performance, enabling models to learn more effectively from reduced data. Additionally, Sucholutsky and Schonlau proposed soft-label dataset distillation, which uses probabilistic labels for distilled data [44], facilitating the training of complex models with minimal data while maintaining strong performance. Despite its success in standard deep learning, dataset distillation is often impractical for robust learning due to its high computational demands. Moreover, it assumes that all the examples in the entire dataset are equally important for distillation, ignoring the fact that different data points may have varying influences on model performance.
Coreset Selection. On the other hand, coreset selection [45, 46] aims to identify a small, representative data subset that retains the model performance when trained on this reduced dataset. These approaches often rely on diversity-based metrics or optimization techniques to ensure that the selected subset preserves sufficient information for effective learning. For example, Kaushal et al. utilized diverse models to select training data subsets to reduce labeling efforts [20], while Xia et al. introduced the concept of “moderate coreset” [21], focusing on data points with scores near the median to construct subsets that generalize well across different scenarios based on score distributions. Similarly, Mirzasoleiman et al. proposed selecting a weighted subset of training data that approximates the full gradient by maximizing a submodular function [22].
In the adversarial context, Dolatabadi et al. proposed adversarial coreset selection [47], offering a task-specific solution to enhance the computational efficiency of adversarial training. In particular, the method constructs a compact data subset by minimizing the gradient approximation error. The constructed subset is then utilized for adversarial training, with selection taking place at regular intervals. While effective in reducing computational overhead for vanilla adversarial training, the growing reliance on large volumes of generated or external data—such as in self-supervised adversarial training—introduces new challenges. These approaches often expand the dataset size to improve robust accuracy, eventually increasing the total training time. In such settings, repeatedly selecting data during training may become computationally infeasible. In summary, our work complements the above-mentioned methods for efficient deep learning, focusing on developing more data selection schemes without compromising robustness or scalability, particularly for SSAT when large-scale unlabeled data are involved.
Appendix B Detailed Experimental Setup
SVHN. SVHN is naturally divided into a core training set comprising approximately 7373K images and an extra training set of around 531531K images. Initially, the model is trained on the 7373K labeled images. To evaluate the generalizability of our selection schemes, we apply it to a synthetic dataset generated using a DDPM model [13].
Configuration. For model training, we adopt a WideResNet 16-8 (WRN-16-8) architecture [48] in all our SVHN experiments. We generate adversarial examples using PGD attacks exactly as implemented by Zhang et al. [9], with step size $0.007$, $10$-step PGD attack iterations, and $\ell_{\infty}$ perturbation magnitude $\epsilon=0.015$. Hyperparameters are set the same as in Carmon et al. [11] except for the number of epochs. For SVHN, we use a training batch size of $128$. All of our experiments that utilize all of the extra 531K data are run for $200$ epochs and the experiments with 11M generated data are run for $400$ epochs. To prevent overfitting and reduce computational complexity, we run our experiments with selected data for $75$ epochs with early stopping. Following prior literature [11], we adopt an SGD optimizer with a weight decay factor of $5\cdot 10^{4}$ and an initial learning rate of $0.1$ with a cosine learning rate annealing to train the model.
Evaluation. For the attack evaluation to calculate the robust accuracy using PGD attack, we keep the parameters similar to that of Carmon et al. [11] for better comparison. We use step size $0.005$, number of attack steps $K=100$. We evaluate models at $\epsilon=0.015$, which is the same as the value we used during training. We also use Auto Attack with the same $\epsilon$ value as in PGD attack.
CIFAR-10. The CIFAR-10 dataset has 50K labeled images. We use the $80$ Million Tiny Images (80M-TI) dataset, of which CIFAR-10 is a manually labeled subset, to obtain extra data. However, most images in 8080M-TI do not correspond to CIFAR-10 image categories. Carmon et al. [11] used an 11-way classifier to distinguish CIFAR-10 classes and an 11th “non-CIFAR-10” class using a WideResNet 28-10 model. Each class selected an additional 5050K images from 80M-TI using the model’s predicted scores to create a 500500K pseudo-labeled dataset, which we used in our experiments. We train the intermediate model with 5050K labeled data and use this model to select data from the 500500K pseudo-labeled data or a synthetic dataset. We perform the same experiments as in SVHN, except that we do not have the ground-truth labels of the additional unlabeled dataset. We also conducted further experiments utilizing 11M synthetical data generated by the DDPM model [13]. For our fine-tuning experiments, we start with a publicly available pretrained DDPM model [49] and further fine-tune it for a few epochs to generate the desired images, which are subsequently used for our final training.
Configuration. For our main CIFAR-10 experiments, we use a WideResNet 28-10 (WRN-28-10) [48] architecture to train models. Similar to SVHN experiments, we set step size $0.007$, PGD attack iterations as $10$ and $\ell_{\infty}$ perturbation magnitude $\epsilon=0.031$. For our adversarial training with $\ell_{2}$ perturbation, we use $\epsilon=128/255$. In section D-A when $\epsilon=4/255$, we use $0.0035$ as the PGD step size. For $\epsilon=16/255$, we use $0.015$ as the PGD step size. Hyperparameters used are the same as in Carmon et al. [11] except for the number of epochs. Here, we use a training batch size of $256$. The experiments using all the 500500K pseudo-labeled data are run for $200$ epochs, and using 1M generated data are run for $400$ epochs. For our experiments with limited data, we get the best results at $100$ epochs where we early stop the training process. Following prior literature [11], we adopt an SGD optimizer with a weight decay factor of $5\cdot 10^{4}$ and an initial learning rate of $0.1$ with a cosine learning rate annealing to train the model.
Evaluation. The PGD attack evaluation is conducted similarly to that of Carmon et al. [11] for fair comparisons. We use step size $\alpha=0.01$, number of attack steps $K=40$. We consider $\ell_{\infty}$ perturbations with $\epsilon=0.031$, the same as the perturbation magnitude we used during training. For the models trained with $\ell_{2}$ perturbation, we evaluate with $\ell_{2}$ perturbation magnitude $\epsilon=128/255$. We also use Auto Attack with the same $\epsilon$ value as in PGD attack.
Medical Application. Below, we detail the setup employed for our experiments on the medical datasets outlined in Section V-A. For this experiment, we trained models using a ResNet-18 [50] architecture We consider $\ell_{\infty}$ perturbations with $\epsilon=0.1$. We set the attack step size to $0.02$, and the number of PGD steps to $10$. The final model is trained with SSAT using the Adam optimizer [51] with a learning rate of $0.001$. We evaluate model robustness with PGD attacks with a step size of $0.02$ and $20$ attack steps.
Appendix C Additional Experiments
(a) CIFAR-10
(b) SVHN
Figure 7: Illustration of clean and robust accuracy curves of SSAT algorithms with LCS-KM by varying per-batch ratio $\gamma$, which defines how much labeled and unlabeled data are included in each mini-batch, on: (a) CIFAR-10 with $20\%$ extra unlabeled data selected from 11M generated images using DDPM, and (b) SVHN with $10\%$ selected from the external 531531K SVHN images.
TABLE IV: Comparison results of semi-supervised adversarial training algorithms with varying data selection schemes on SVHN. Instead of pseudo labels, the selected unlabeled data are assigned the ground-truth SVHN labels provided in the dataset.
$\textbf{Ratio }\bm{\alpha}$ Method Clean Accuracy ($\%$) Robust Accuracy ($\%$) Number of Epochs
$1\%$ LCS-KM $95.6$ $82.9$ $75$
$10\%$ Random $96.3$ $82.7$ $75$
$10\%$ LCS-KM $96.7$ $86.3$ $75$
$20\%$ LCS-KM $96.4$ $86.5$ $75$
$100\%$ No Selection $97.5$ $86.4$ $200$
C-A Hyperparameter Sensitivity Analysis Details
In Section V-C, we present a detailed analysis of how accuracies are influenced by different hyperparameter values. In these experiments, we varied one hyperparameter at a time while keeping the others constant. Figure 4a illustrates how robust accuracy on the CIFAR-10 dataset changes with $\beta$ values when $20\%$ additional data is incorporated from the 11M data samples generated by the DDPM model, using a per-batch ratio of $0.3$. Figure 4b shows the variation in robust accuracy for the SVHN dataset as $\beta$ changes, with $10\%$ drawn from the SVHN extra dataset and a per-batch ratio of $0.5$.
In Figure 7a, we present how SSAT’s robust accuracy is affected on CIFAR-10 by varying $\gamma$ when $20\%$ extra data from the 11M generated DDPM samples is used, with $\beta$ fixed at $0.4$. Similarly, Figure 7b displays the effect of varying the per-batch ratio on robust accuracy for the SVHN dataset, using $10\%$ data from the SVHN extra dataset and fixing $\beta$ at $0.6$. In addition, Figures 4c and 4d explore how robust accuracy varies with the amount of data used. For CIFAR-10 experiments in Figure 4c, the per-batch ratio is set to $0.5$, and $\beta$ is fixed at $0.4$. For SVHN in Figure 4d, the per-batch ratio is set to $0.5$, and $\beta$ is set to $0.6$. We aim to capture the robust accuracy at the optimal epoch in each case, and we observe that for Figures 4c and 4d, as the amount of data increases, the optimal accuracy is reached at a considerably later epoch.
(a) CIFAR-10
(b) SVHN
Figure 8: Comparisons of SSAT methods with varying amounts of unlabeled data selected using LCS-KM with the CLIP model on (a) CIFAR-10 and (b) SVHN. We report standard and robust accuracies on the labeled test dataset.
C-B Pre-Trained CLIP as Intermediate Model
In our experiments, we always began by training an intermediate model $f_{\hat{\theta}}$ using the same architecture as the final model $\theta_{\mathrm{final}}$. $f_{\hat{\theta}}$ was trained exclusively on the labeled dataset and then used to generate pseudo labels and select a subset from the extra dataset. However, when dealing with a generated or external dataset that already comes with labels, training an intermediate model specifically for selection undermines the data selection process’s intended benefits, primarily to reduce computational and time complexity. In application scenarios where the size of the labeled dataset $\mathcal{S}_{l}$ is also limited, likely, the intermediate model can not be trained with high accuracy, leading to large amounts of inaccurate pseudo labels.
To address these issues, we study the feasibility of using an alternative pre-trained model in LCS to perform data selection, specifically leveraging the CLIP model for extracting image embeddings [52]. CLIP, trained on a vast array of internet images and text, offers a rich and versatile feature space that can be utilized without additional training. By using the capabilities of pre-trained CLIP, we can bypass the intermediate step of training a model from scratch, thereby streamlining the selection process and further enhancing overall efficiency. This substitution allows us to select a relevant subset without having to train an intermediate model from scratch. Our experiments indicate that using CLIP for data selection yields results that are comparable to those obtained using the trained intermediate model, both in terms of standard and robust accuracies. Figure 8 shows our result when using CLIP instead of the previous intermediate model for data selection.
C-C SVHN with Ground-Truth Labels
We conduct additional experiments to test the performance of our data selection scheme on SVHN with pseudo labels replaced by ground-truth labels for the extra data. Table IV shows the comparison results. With ground-truth labels, our LCS schemes can achieve similar robust accuracy by selecting just $10\%$ extra data compared with no selection result.
C-D Resnet-18 Model Architecture
Our main experiments focus on the WideResNet architecture. We further evaluate the performance of SSAT under $\ell_{\infty}$ perturbations with $\epsilon=0.031$ using different data selection schemes and varying selection ratios with external unlabeled data on CIFAR-10 using a ResNet-18 model. Similar to our previous experiments, we train the ResNet-18 model on 5050K labeled samples from CIFAR-10 and a chosen subset of data from 500500K unlabeled images of Tiny-ImageNet. Our evaluation results are shown in Table V, which suggests that our proposed LCS methods can greatly improve the efficacy of SSAT on ResNet-18 while achieving comparable robustness performance of “No Selection” with $\alpha=100\%$.
TABLE V: Comparisons of SSAT performance on CIFAR-10 using the ResNet-18 architecture under various configurations. The model training dataset consists of 5050K labeled CIFAR-10 images and 500500K unlabeled images drawn from Tiny-ImageNet.
$\textbf{Ratio }\bm{\alpha}$ Method Clean ($\%$) PGD ($\%$)
$1\%$ Random $79.8$ $50.6$
PCS $79.4$ $51.1$
LCS-GMM $78.7$ $51.6$
LCS-KM $78.1$ $51.8$
$10\%$ Random $82.3$ $52.0$
PCS $82.1$ $52.6$
LCS-GMM $82.0$ $53.2$
LCS-KM $83.3$ $53.6$
$20\%$ Random $82.4$ $53.1$
PCS $82.1$ $53.3$
LCS-GMM $82.1$ $53.7$
LCS-KM $83.1$ $54.0$
$100\%$ No Selection $83.2$ $54.7$
Appendix D Extended Analyses
In this section, we further evaluate the flexibility of our method across various adversarial training algorithms and perturbation configurations (Section D-A), and examine the impact of intermediate model quality, particularly under the regime with insufficient labeled data (Section D-B). Without explicit mention, we use our LCS-KM method for selecting critical unlabeled data, as it derives the best-performing model when integrated with SSAT as suggested by Table I.
(a) TRADES ($\epsilon=4/255$, PGD-10)
(b) TRADES ($\epsilon=16/255$, PGD-10)
(c) AT ($\epsilon=8/255$, PGD-5)
(d) AT ($\epsilon=8/255$, PGD-10)
Figure 9: Illustration of SSAT performance on CIFAR-10 using LCS-KM under $\ell_{\infty}$ perturbations under various configurations: TRADES using $10$ PGD steps with (a) $\epsilon=4/255$ and (b) $\epsilon=16/255$; AT with $\epsilon=8/255$ using (c) 55 PGD steps and (d) $10$ PGD steps.
TABLE VI: Performance of intermediate models trained using different schemes, including self-supervised learning (SSL), pre-trained CLIP, and fully-supervised methods, on CIFAR-10 under varying ratios of labeled data. We also present the corresponding performance of the final model produced by SSAT with LCS-KM, where the ratio of selected unlabeled data $\alpha$ is set at $10\%$.
Labeled Ratio Intermediate Model Final Model (LCS-KM, $\alpha=10\%$)
Training Method Clean Acc ($\%$) Clean Acc ($\%$) Robust Acc ($\%$)
$1\%$ SSL ($5\%$ unlabeled) $78.8$ $82.9$ $39.8$
pre-trained CLIP $86.8$ $83.9$ $40.5$
$10\%$ SSL ($5\%$ unlabeled) $82.9$ $85.0$ $42.4$
pre-trained CLIP $86.8$ $85.3$ $43.2$
$20\%$ SSL ($5\%$ unlabeled) $87.0$ $87.2$ $48.3$
pre-trained CLIP $86.8$ $85.6$ $47.7$
100% fully-supervised $90.5$ $86.2$ $58.2$
pre-trained CLIP $86.8$ $86.5$ $57.4$
D-A Further Exploration of Generalizability
We compare four different settings of SSAT on CIFAR-10: random selection with $\alpha$ ratio of unlabeled data, selecting $\alpha$ ratio using LCS-KM, using only labeled data ($\alpha=0\%$), and using the entire dataset ($\alpha=100\%$).
Perturbation Size. We consider external data from TinyImages as the unlabeled dataset. We investigate the effect of varying the $\ell_{\infty}$ perturbation size $\epsilon$, setting it to $4/255$ and $16/255$ while keeping TRADES as the adversarial training algorithm similar to the experiments in Table I. The results are depicted in Figures 9a and 9b, suggesting that our LCS-KM method is highly effective compared with random selection and no selection, almost approaching the robust accuracies achieved by full utilization of unlabeled data. The trends are shown to be consistent across varying perturbation sizes.
Adversarial Training. We explore the impact of replacing TRADES with vanilla AT as the training algorithm. Similarly, we use external data from TinyImages as the unlabeled dataset. The results for PGD-based AT are shown in Figure 9d. Additionally, we reduce the number of PGD steps from $10$ to $5$, and the outcomes are summarized in Figure 9c. Across all configurations, our LCS-KM selection consistently outperforms random selection and achieves accuracy comparable to using the entire labeled dataset, suggesting the generalizability and robustness of our methods.
(a) External data
(b) Generated data
Figure 10: Illustration of SSAT performance on CIFAR-10 under $\ell_{2}$ perturbations with $\epsilon=128/255$ using LCS-KM by varying $\alpha$: (a) external TinyImages and (b) DDPM-generated data.
$\bm{\ell}_{2}$ Perturbations. We further evaluate the performance of LCS-KM under $\ell_{2}$ perturbations with size $\epsilon=128/255$. In this experiment, we augment CIFAR-10 with either external unlabeled data sourced from Tiny-ImageNet or synthetic data generated via DDPM. We consider three configurations of selecting $10\%$, $20\%$, and $100\%$ extra unlabeled data. Our findings are shown in 10a and 10b, revealing that selecting $20\%$ unlabeled data achieves nearly equivalent robust accuracy compared to using the full dataset with $\alpha=100\%$. These results again confirm the efficiency and effectiveness of LCS-KM when adopted in SSAT algorithms.
D-B Low Labeled Data Regime
We further study how the choice of the intermediate model and the amount of labeled data impact the final performance. For all experiments, we use a subset of the CIFAR-10 dataset as labeled data, supplemented with $10\%$ of data selected from a pool of 500500K unlabeled images sourced from Tiny-ImageNet. This additional data is selected using the intermediate model. The amount of unlabeled data remains constant across experiments, while the quantity of labeled data is varied. The intermediate model is trained using self-supervised learning (SSL) with a $5\%$ randomly selected subset of the unlabeled dataset, as long as the labeled data ratio is low. In addition, we include a pre-trained CLIP model [52] as an alternative intermediate model for comparison. The results of these experiments are presented in Table VI. The intermediate model’s quality significantly impacts the final model’s robust accuracy. In low-label regimes (e.g., $1\%$ or $10\%$), models trained using our intermediate model yield comparatively lower robust accuracy than those leveraging the pre-trained CLIP model. However, as the amount of labeled data for the intermediate model increases (e.g., 20% or 100%), the final model results are similar to or better than those of its counterparts based on the CLIP intermediate model. This trend demonstrates the effect of the intermediate model on the final robust accuracy. These observations underscore the importance of having a strong intermediate model for robust performance. Besides, we observe that pre-trained models like CLIP can be a good alternative to training an intermediate model, especially in low-labeled-data regimes. Detailed experiment results and discussions on the pre-trained CLIP intermediate model are provided in Appendix C-B.
D-C Further Discussion
Our work introduces several data reduction strategies that tackle the key challenges in adversarially robust learning, particularly the significant memory and computational overhead of state-of-the-art SSAT methods. By prioritizing boundary-adjacent data, these processes aim to enhance model robustness. However, these methods assume that the model’s latent representations are reliable for clustering and confidence evaluation, an assumption that may falter in the case of undertrained or poorly calibrated models. While these approaches effectively reduce the volume of unlabeled data required for training, their performance can be sensitive to hyperparameter configurations, such as the balance between boundary and non-boundary points and the size of the unlabeled dataset. By focusing on challenging examples near the decision boundary, these selection methods naturally address some aspects of robust overfitting and support robust generalization. Nevertheless, a deeper investigation is required to fully comprehend how these strategies influence decision boundary geometry and the model’s capacity for robust generalization. Combining these selection techniques with fast and efficient adversarial example generation methods, such as fast and free adversarial training, holds promise for further accelerating the training process while preserving robust performance.
Reducing the overall time required for adversarial training is critical for improving the efficiency and scalability of robust learning. Broadly, there are two complementary strategies to achieve this: (i) decreasing the computational cost of each training epoch, and (ii) reducing the total number of epochs needed for convergence. The first strategy aims to make each epoch faster without sacrificing robustness. Several prior works [18, 19, 53, 54] focus on this direction by introducing more efficient attack generation and optimization schemes. These approaches demonstrate that careful algorithmic design can substantially reduce per-epoch runtime, making adversarial training far more practical in large-scale settings. However, as shown in previous studies [15, 13], the total number of training epochs grows sharply with the size of the dataset. Consequently, even if the cost per epoch is reduced, increasing the amount of training data can still lead to a dramatic rise in overall training time. For example, adversarial training with approximately 11M generated samples typically requires around $400$ epochs to converge, whereas training with 100100M samples may demand as many as $2000$ epochs. Developing strategies that mitigate this dependence so that large-scale training does not necessitate a proportional increase in epochs would provide substantial computational savings and further improve the practicality of adversarially robust learning.